SPECTROSCOPY OF SUBSTITUTED CYCLOHEXOXY RADICALS AND KINETICS OF THEIR REAC-TIONS WITH O_{2}

\underline{THEODORE S. DIBBLE}, LEI ZHANG, Chemistry Department, SUNY-Environmental Science and Forestry, Syracuse, NY 13210; KAREN M. CALLAHAN, Chemistry Department, The Ohio State University, 120 W. 18th Ave., Columbus, OH 43210; and DEAN A. DERBYSHIRE, Chemistry Department, Le Moyne College, Syracuse, NY, 13214.

The laser induced fluorescence spectrum of the B̃ state of 4-methyl cyclohexoxy radical was obtained at 228 K and 50 Torr N_{2}. The spectrum appears to arise only from the trans isomer, and is very similar to that observed from cyclohexoxy radical, itself. Thermodynamic arguments indicate the spectrum must arise almost solely from the di-equatorial chair conformer. No significant fluorescence was observed from 4-tert-butyl cyclohexoxy radical. Rate constants were obtained for the reaction of trans-4-methyl cyclohexoxy radical and d11-cyclohexoxy radical with O_{2}. Rate constants were measured between 228 and 301 K, and are independent of pressure (50-125 Torr). The temperature dependence of the rate constant for trans-4-methyl cyclohexoxy radical is not as great as that previously measured for cyclohexoxy radical, but is greater than that found previously for acyclic alkoxy radicals.